# An Adverse Weather-Immune Scheme with Unfolded Regularization and Foundation Model Knowledge Distillation for Street Scene Understanding

### v1

@@ 8                   $PF_{t}\leftarrow\text{CNN}_{\text{shd}}(MF_{t})$
12                   $PInsU_{t}\leftarrow\text{CNN}_{\text{InsU}}(InsU_{t})$
13                   $PIntU_{t}\leftarrow\text{CNN}_{\text{IntU}}(IntU_{t})$
14                   $PDU_{t}\leftarrow\text{CNN}_{\text{DU}}(DU_{t})$
15                   $PF_{t}\leftarrow\text{Merge}(PInsU_{t},PIntU_{t},PDU_{t})$
@@ 20       $MiniBatches\leftarrow\text{Shuffle\&Group }\{PF_{t}\}$
22             $\text{Head},\text{CNN}_{\text{shd}}\ (or\ \text{CNN}_{\text{InsU/IntU/DU}})\leftarrow B$
@@ III-B Unfolded Regularization
On top of the conventional pixel-level cross entropy loss (denoted as $L_{CE}$) in semantic understanding, we propose an image-level regularizer and an inter-class contrastive regularizer to enhance the model’s across-weather generalization. URs are the deep unfolding implementation of such regularizers, which aims at avoiding the heuristics and exhaustive searches on weights of them.
III-B1 The proposed Regularizers
We denote the distrituion of predicted masks $X$ and the ground truth $Y$ as $P_{X}$ and $P_{Y}$, respectively. We propose using Bhattacharyya distance (BD) as the image-level regularizer to quantify the distribution similarity between $X$ and $Y$. The definition of the BD regularizer over pixel $x$ is given by:
$L_{BD}=D_{B}(P_{X},P_{Y})=-\ln(\sum\nolimits_{x}\sqrt{P_{X}(x)P_{Y}(x)}),$ (1)
On the other hand, we propose using InfoNCE as an inter-class contrastive regularizer that can pull pixles of same class closer while push pixles of different classes apart. The InfoNCE regularizer is expressed as:
@@ III-B1 The proposed Regularizers
where $x$ is an anchor pixel, $u_{m}$ is any pixel from the same-class set $\mathcal{U}$, $v_{n}$ is any pixel from the different-class set $\mathcal{V}$, $\sigma(\cdot,\cdot)$ denotes the inner product, and $\tau$ controls the separation sharpness.
Figure 3: Illustration of URs. (4), (5), (6) represent equation number.
@@ III-B2 Unfolded Regularizers (URs)
The total loss $L$ for training is:
$L=L_{CE}+\alpha L_{BD}+\gamma L_{con},$ (3)
where $\alpha$ and $\gamma$ are the weights assigned to $L_{BD}$ and $L_{con}$, respectively. In practice, they are tuned based on heuristics or exhaustive searches, which results in instability and suboptimal performance. Treating $\alpha$ and $\beta$ as trainable parameters can avoid these issues.
To this end, inspired by [7], we propose to treat each optimization iteration as one layer, and treat $\alpha$ and $\beta$ as learnable parameters within each layer. Based on this, above both regularizers can be unfolded into $K$ layers (illustrated in Fig. 3). Specifically, for layer $k,\text{where}\ k\in\{1,2,\cdots,K\}$, the output is updated as follow:
$x^{(k)}=\hskip 4.26773pt$ $x^{(k-1)}-\eta^{(k)}(\alpha^{(k)}\nabla_{x}L_{BD}(x^{(k-1)})\hskip 4.26773pt+\hskip 4.26773pt$
$\gamma^{(k)}\nabla_{x}L_{con}(x^{(k-1)})),$ (4)
where $\alpha^{(k)}$ and $\gamma^{(k)}$ are the learnable weights, $\eta^{(k)}$ is the step size of update, $\nabla_{x}L_{BD}(x^{(k-1)})$ and $\nabla_{x}L_{con}(x^{(k-1)})$ are the gradients of $L_{BD}$ and $L_{con}$ relative to pixel $x^{(k-1)}$, respectively, and they can be formulated as follows (denoting $x^{(k-1)}$ as $x$ for short):
@@
After multiple gradient descent updates by $K$ sequentially-connected layers, the output of the $K$-th layer is optimized by both regurlarizers, and then is added to the output of LSaGS. Then we use the summation results serving as predicted mask and the ground truth to calculate the cross entropy loss to optimize LSaGS and $\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}$ by back propagation. In conclusion, URs is outlined in Algo. 2.
Input: $X$ (Input image), $Y$ (One-hot ground truth), $K$ (Layer number)
Output: Model $LSaGS$, Learnable variables $\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}$
1 Initialize model $LSaGS\leftarrow\mathcal{W}_{0}$, and $\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)}\leftarrow\alpha_{0},\gamma_{0},\eta_{0},\text{where}\ k\in\{1,2,\cdots K\}$
@@
3       $x^{(0)}\leftarrow LSaGS(X)$
       // Unfolding layers
@@ 3       $x^{(0)}\leftarrow LSaGS(X)$
10      $x\leftarrow x^{(0)}+x^{(K)}$
11       $L=CrossEntropy(x,Y)$
12       $LSaGS,\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}\leftarrow L.Backward()$
@@ 11       $L=CrossEntropy(x,Y)$
Algorithm 2 Unfolded Regularizers (URs)
III-C SAM-Aided Consecutive Frame Annotation
SAM, a vision foundation model known for its strong zero-shot inference abilities, is used to generate masks for images in our research. It should be noted that SAM operates online during the training phase and offline during the inference phase, as illustrated in Fig. 4. However, SAM does not support generating category information (as demonstrated in LABEL:Fig.SAM_reason_to_cluster). To overcome this limitation, SBICAC algorithm is introduced to assign each semantic class a unique ID, which are then served as labels for training the model. SBICAC follows an iterative refinement manner and each iteration is described as follows:
@@ III-C SAM-Aided Consecutive Frame Annotation
In summary, SBICAC leverages inner product similarity to iteratively assign pixels to clusters and update clusters’ centroids until the cluster assignments stabilize. LABEL:Fig.cluster_comp exemplifies the performance of SBICAC against other clustering algorithms. This method ensures that the clusters are formed based on maximizing the similarity between pixels and their respective centroids. Upon completion of the iterative process, SBICAC outputs the cluster assignments $labels$ for all pixels. SBICAC is outlined in Algo. 3.
@@
1 Initialize centroids as $centroids\leftarrow init$
@@
3       Compute similarity between $X$ and $centroids$: $similarity\leftarrow X\cdot centroids^{T}$
4       Assign clusters based on maximum similarity: $labels\leftarrow\arg\max(similarity,axis=1)$
@@ 5       Cluster Centroid Calculation:
7             $cluster\ j$: $new\_centroids[j]\leftarrow\frac{1}{|X_{j}|}\sum X_{j}$
@@ 9      Check for convergence:
14      Update centroids: $centroids\leftarrow new\_centroids$
@@ IV-A1 Datasets
In our research, we utilize two datasets: the Apolloscapes dataset [30] and the CARLA_ADV dataset derived from the Carla simulator (version 0.9.13) [31]. The Apolloscape dataset is a comprehensive AD dataset featuring a variety of driving scenarios captured under various weather conditions. Each annotated image has pixel-level labels across 23 classes such as vehicles and pedestrians. The CARLA_ADV dataset is specifically designed to encompass a range of adverse weather conditions including fog, clouds, rain, darkness, and combinations thereof. It does also contain 23 semantic classes in pixel-level, including vehicle, building, tree, etc.
@@ IV-A2 Metrics
We evaluate the proposed AdvImmu on street scene understanding task by employing four metrics: Mean Intersection over Union (mIoU), which measures the overlap between predicted mask and ground truth; Mean Precision (mPre), which assesses the accuracy of positive predictions; Mean Recall (mRec), evaluating how well the model identifies all relevant instances; and Mean F1 (mF1), which provides a balance between precision and recall.
TABLE I: The performance comparison of AdvImmu against other baselines
Methods Apolloscape CARLA_ADV
@@
DeepLabv3+ [32] 26.58$\pm$0.49 30.23$\pm$0.89 31.14$\pm$0.57 32.32$\pm$0.59 36.90$\pm$1.00 46.96$\pm$0.97 42.44$\pm$0.79 43.78$\pm$0.93
BiSeNetV2 [33] 22.92$\pm$0.86 27.85$\pm$1.20 27.10$\pm$1.11 27.12$\pm$1.57 28.80$\pm$1.93 34.47$\pm$2.23 33.46$\pm$2.19 33.59$\pm$2.30
SegNet [34] 21.01$\pm$0.51 24.80$\pm$1.11 24.96$\pm$0.54 24.60$\pm$0.70 31.67$\pm$2.11 38.45$\pm$3.09 36.55$\pm$2.44 37.15$\pm$2.70

### v3

@@ 8                   $PF_{t}\leftarrow\text{CNN}_{\text{shd}}(MF_{t})$;
12                   $PInsU_{t}\leftarrow\text{CNN}_{\text{InsU}}(InsU_{t})$;
13                   $PIntU_{t}\leftarrow\text{CNN}_{\text{IntU}}(IntU_{t})$;
14                   $PDU_{t}\leftarrow\text{CNN}_{\text{DU}}(DU_{t})$;
15                   $PF_{t}\leftarrow\text{Merge}(PInsU_{t},PIntU_{t},PDU_{t})$;
@@ 20       $MiniBatches\leftarrow\text{Shuffle\&Group }\{PF_{t}\}$
22             $\text{CNN}_{\text{shd}}\ (or\ \text{CNN}_{\text{InsU/IntU/DU}})\leftarrow B$;
@@ III-B Unfolded Regularization
On top of the conventional pixel-level cross entropy loss (denoted as $L_{CE}$) in semantic understanding, we propose an image-level regularizer and an inter-class contrastive regularizer to enhance the model’s across-weather generalization. URs are the deep unfolding implementation of such regularizers, which aims at avoiding the heuristic and exhaustive searches on weights of them.
III-B1 Definition of the proposed Regularizers
To formulate the definition clearly, we denote each input image as $F$, the LSaGS’s predicted masks as $X$ and the corresponding ground truth as $Y$. Furthermore, we term each pixel’s prediction within $X$ as $x$, and denote the distribution of $X$ and $Y$ as $P_{X}$ and $P_{Y}$, respectively.
On the one hand, we propose to use Bhattacharyya distance (BD) as the image-level regularizer to quantify the distribution similarity between $P_{X}$ and $P_{Y}$. The definition of the BD regularizer over input image $X$ is given by:
$L_{BD}=D_{B}(P_{X},P_{Y})=-\ln(\sum\nolimits_{x\in X}\sqrt{P_{X}(x)P_{Y}(x)}),$ (1)
On the other hand, we propose to use InfoNCE as an inter-class contrastive regularizer that can pull pixles of same class closer while push pixles of different classes apart. The InfoNCE regularizer is expressed as:
@@ III-B1 Definition of the proposed Regularizers
where $x$ is an anchor pixel, $u_{m}$ is any pixel from the same-class pixle set $\mathcal{U}$, $v_{n}$ is any pixel from the different-class pixel set $\mathcal{V}$, $\sigma(\cdot,\cdot)$ denotes the inner product, and $\tau$ controls the separation sharpness.
Figure 5: Illustration of AdvImmu architecture, which contains LSaGS and URs. (4), (5), (6) represent equation number.
@@ III-B2 Unfolded Regularizers (URs)
On top of the conventional cross entropy loss $L_{CE}$, considering the proposed $L_{BD}$ and $L_{con}$, the total loss $L$ for semantic segmentation task is
$L=L_{CE}+\alpha L_{BD}+\beta L_{con},$ (3)
where $\alpha$ and $\beta$ are the weights assigned to $L_{BD}$ and $L_{con}$, respectively. Traditionally, $\alpha$ and $\beta$ are tuned based on heuristic or exhaustive searches, and then the total loss $L$ is used to optimize the model in training phase. This conventional optimization strategy generally results in instability and suboptimal performance because there is no guarantee optimal $\alpha$ and $\gamma$ by a finite number of searches. To tackle this issue, we propose a two-stage optimization policy based on above discussed $L_{BD}$ and $L_{con}$ regularizers.
For the first stage, we propose to convert weights $\alpha$ and $\beta$ as trainable parameters rather than searched hyper-parameters. To this end, inspired by [18, 70], we propose to treat each optimization iteration of the proposed $L_{BD}$ and $L_{con}$ as one layer, and treat $\alpha$ and $\beta$ as learnable parameters within each layer. Based on this, both regularizers can be unfolded into $K$ layers (illustrated in the right part of Fig. 5) by using deep unfolding technique. Then such $K$ unfolded layers are appended at the end of the LSaGS model, serving as the part of AdvImmu model architecture.
To formulate the update rule of URs, for layer $k,\text{where}\ k\in\{1,2,\cdots,K\}$, the output $x^{(k)}$ corresponding to pixel $x\in X$ is updated as follow:
$x^{(k)}=\hskip 4.26773pt$ $x^{(k-1)}-\eta^{(k)}(\alpha^{(k)}\nabla_{x^{(k-1)}}L_{BD}(x^{(k-1)})\hskip 4.26773pt+\hskip 4.26773pt$
$\beta^{(k)}\nabla_{x^{(k-1)}}L_{con}(x^{(k-1)})),$ (4)
where $x^{(0)}$ is the input of URs and it exactly equals to $x\in X$, $\alpha^{(k)}$ and $\beta^{(k)}$ are the learnable weights, $\eta^{(k)}$ is the learnable step size of update, $\nabla_{x^{(k-1)}}L_{BD}(x^{(k-1)})$ and $\nabla_{x^{(k-1)}}L_{con}(x^{(k-1)})$ are the gradients of $L_{BD}$ and $L_{con}$ relative to pixel $x^{(k-1)}$, respectively, and they can be formulated as follows (denoting $x^{(k-1)}$ as $x$ for short):
@@
After multiple gradient descent updates by $K$ sequentially-connected layers, the output of the $K$-th layer is optimized by both regurlarizers.
For the second stage, the optimized output of the $K$-th unfolded layer is added to the output of LSaGS model. Then we use the summation results serving as predicted masks and the ground truth to calculate the cross entropy loss $L_{CE}$ to optimize LSaGS and $\alpha^{(k)}$, $\beta^{(k)}$, $\eta^{(k)}\},\text{where}\ k\in\{1,2,\cdots K\}$ by back propagation. Notably, this stage follows the traditional model optimization.
In conclusion, the proposed AdvImmu contains LSaGS and URs, which are collectively illustrated in Fig. 5. In addition, AdvImmu is further outlined in Algorithm 2, where $K$ is a hyper-parameter and can be determined according to practical needs and conditions.
Input: $F$ (Input image from sequence), $Y$ (One-hot ground truth), $K$ (Layer number)
Output: Model $LSaGS$, Learnable variables $\alpha^{(k)}$, $\beta^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}$
1 Initialize model $LSaGS\leftarrow\mathcal{W}_{0}$, and $\alpha^{(k)}$, $\beta^{(k)}$, $\eta^{(k)}\leftarrow\alpha_{0},\gamma_{0},\eta_{0},\text{where}\ k\in\{1,2,\cdots K\}$;
@@
       // Forward propagation
3       $x^{(0)}\leftarrow LSaGS(F)$;
@@
10      $\hat{x}\leftarrow x^{(0)}+x^{(K)}$;
       // Back propagation
11       $L_{CE}=CrossEntropy(\hat{x},Y)$;
12       $LSaGS,\alpha^{(k)}$, $\beta^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}\leftarrow L_{CE}.Backward()$;
@@ 11       $L_{CE}=CrossEntropy(\hat{x},Y)$;
Algorithm 2 AdvImmu
III-C SBICAC’s Annotation based on SAM-distilled Knowledge
The training of the proposed AdvImmu requires frame-wise annotation for consecutive frames, which is impractical to be done manually. To generate the required frame-wise annotation automatically, we propose an unsupervised clustering algorithm based on the distilled knowledge from a large vision model SAM.
SAM, a vision foundation model known for its strong zero-shot inference abilities, is used to assist generating masks for images in our research. It should be noted that SAM operates online during the training phase and offline during the inference phase, as illustrated in the dashed box in Fig. 6. However, SAM does not support generating category information (as demonstrated in LABEL:Fig.SAM_reason_to_cluster). To overcome this limitation, SBICAC algorithm is introduced to assign each semantic class a unique ID based on SAM-distilled latent features, which are then served as labels for training the model. SBICAC follows an iterative refinement manner and each iteration is described as follows:
@@ III-C SBICAC’s Annotation based on SAM-distilled Knowledge
In summary, SBICAC leverages inner product similarity to iteratively assign pixels to clusters and update clusters’ centroids until the cluster assignments stabilize. SBICAC ensures that the clusters are formed based on maximizing the similarity between pixels and their respective centroids. Upon completion of the iterative process, SBICAC outputs the cluster assignments $labels$ for all pixels. The whole pipeline of using SBICAC+SAM to generate frame-wise annotation is illustrated in Fig. 6, and SBICAC clustering algorithm is outlined in Algorithm 3.
In addition, LABEL:Fig.cluster_comp exemplifies the annotation of SBICAC against other off-the-shelf clustering algorithms for a specific image. We can observe that KMeans exists the issue of small object error, while AgglomerativeClustering and SpectralClustering face the problem of larger boundary uncertainty. Obviously, the proposed SBICAC achieves relatively better performance compared to those off-the-shelf competitors.
@@
1 Initialize centroids as $centroids\leftarrow init$;
@@
3       Compute similarity between $X$ and $centroids$: $similarity\leftarrow X\cdot centroids^{T}$;
4       Assign clusters based on maximum similarity: $labels\leftarrow\arg\max(similarity,axis=1)$;
@@ 5       Cluster Centroid Calculation:
7             $cluster\ j$: $new\_centroids[j]\leftarrow\frac{1}{|X_{j}|}\sum_{k=1}^{|X_{j}|}X_{j}^{(k)}$;
@@ 9      Check for convergence:
14      Update centroids: $centroids\leftarrow new\_centroids$;
@@ IV Experiments
In this section, we firstly introduce the experimental settings in Section IV-A. Based on such settings, we will evaluate the proposed AdvImmu’s performance, convergence, complexity and inference delay in Section IV-B. Subsequently, we conduct ablation studies about AdvImmu in Section IV-C. Finally, we will evaluate SBICAC in Section IV-D.
@@ IV-A1 Datasets
In the evaluation, we utilize six datasets:
  • The Apolloscapes dataset [71] is a comprehensive AD dataset featuring a variety of driving scenarios captured under various weather conditions. For our purposes, we use a subset of this dataset, which includes 854 training images and 400 test images, each annotated with pixel-level labels across 23 classes such as vehicle, pedestrian, etc.
  • The CARLA_ADV dataset derives from the Carla simulator (version 0.9.13) [72]. CARLA_ADV dataset is specifically designed to encompass a range of adverse weather conditions including fog, clouds, rain, darkness, and combinations thereof. It contains 2,764 training images and 1,921 test images, annotated at the pixel level for 23 classes, including vehicle, building, tree, etc.
  • The Cityscapes dataset [73], captured across multiple cities, comprises 2,975 training images and 500 test images. The dataset features pixel-level labels for 20 classes, such as vehicles, pedestrians, etc. Notably, we disable LSM in AdvImmu in case of the test on Cityscapes dataset because Cityscapes dataset does not contain consecutive frames.
  • The GTA5 dataset [74] comprises 24,966 synthetic images, each with pixel-level semantic annotations. These images are rendered from the video game Grand Theft Auto 5, capturing scenes from ego-vehicle’s perspective in American-style virtual cities. We use a subset of this dataset, including 2,000 training images and 500 test images. The dataset includes 20 semantic classes compatible with those in the Cityscapes dataset.
  • The Shift dataset [75] consists of images captured across various adverse weather conditions and different time slot of the day. This dataset is particularly valuable for developing and testing algorithms that must perform consistently in diverse and dynamically changing real-world settings. The training and test datasets contain 7,218 and 2,875 images, respectively, where each image is annotated with pixel-level labels for 23 semantic classes.
  • The SynthiaSF dataset [76] features a collection of synthetic but photorealistic images that simulate urban driving scenarios. In our experiments, the training dataset and the test dataset include 1,596 and 628 images, respectively. This dataset includes pixel-level annotations for 23 semantic classes.
@@ IV-A2 Metrics
We evaluate the proposed AdvImmu on street scene understanding task by employing four metrics: Mean Intersection over Union (mIoU), which measures the overlap between predicted mask and ground truth; Mean Precision (mPre), which assesses the accuracy of positive predictions; Mean Recall (mRec), evaluating how well the model identifies all relevant instances; and Mean F1 (mF1), which provides a balance between precision and recall. These metrics are formulated as follows:
$mIoU=\frac{1}{\mathcal{C}}\sum_{c=1}^{\mathcal{C}}IoU_{c}=\frac{1}{\mathcal{C}\mathcal{N}}\sum_{c=1}^{\mathcal{C}}\sum_{n=1}^{\mathcal{N}}\frac{TP_{n,c}}{FP_{n,c}+TP_{n,c}+FN_{n,c}},$
$mPre=\frac{1}{\mathcal{C}}\sum_{c=1}^{\mathcal{C}}Pre_{c}=\frac{1}{\mathcal{C}\mathcal{N}}\sum_{c=1}^{\mathcal{C}}\sum_{n=1}^{\mathcal{N}}\frac{TP_{n,c}}{FP_{n,c}+TP_{n,c}},$
$mRec=\frac{1}{\mathcal{C}}\sum_{c=1}^{\mathcal{C}}Rec_{c}=\frac{1}{\mathcal{C}\mathcal{N}}\sum_{c=1}^{\mathcal{C}}\sum_{n=1}^{\mathcal{N}}\frac{TP_{n,c}}{TP_{n,c}+FN_{n,c}},$
$mF1=\frac{1}{\mathcal{C}}\sum_{c=1}^{\mathcal{C}}F1_{c}=\frac{1}{\mathcal{C}}\sum_{c=1}^{\mathcal{C}}\frac{2*Pre_{c}*Rec_{c}}{Pre_{c}+Rec_{c}},$ (11)
where $TP$, $FP$, $TN$ and $FN$ stand for True Positive, False Positive, True Negative and False Negative, respectively. $\mathcal{C}$ denotes the number of semantic classes within the dataset and $\mathcal{N}$ signifies the size of the test dataset. For example, with respect to SynthiaSF dataset, $\mathcal{C}$ is 23 and $\mathcal{N}$ is 628.
TABLE I: The performance comparison of AdvImmu against other baselines across multiple datasets
Methods Apolloscape Dataset CARLA_ADV Dataset
@@
DeepLabv3+ 26.58$\pm$0.49 30.23$\pm$0.89 31.14$\pm$0.57 32.32$\pm$0.59 36.90$\pm$1.00 46.96$\pm$0.97 42.44$\pm$0.79 43.78$\pm$0.93
BiSeNetV2 22.92$\pm$0.86 27.85$\pm$1.20 27.10$\pm$1.11 27.12$\pm$1.57 28.80$\pm$1.93 34.47$\pm$2.23 33.46$\pm$2.19 33.59$\pm$2.30
SegNet 21.01$\pm$0.51 24.80$\pm$1.11 24.96$\pm$0.54 24.60$\pm$0.70 31.67$\pm$2.11 38.45$\pm$3.09 36.55$\pm$2.44 37.15$\pm$2.70
AttaNet 20.59$\pm$0.18 25.88$\pm$0.31 25.38$\pm$0.22 25.55$\pm$0.19 27.12$\pm$1.40 33.41$\pm$1.68 32.90$\pm$1.65 32.60$\pm$1.57
BASeg 20.14$\pm$0.18 31.72$\pm$0.29 25.26$\pm$0.20 25.75$\pm$0.31 36.36$\pm$0.76 57.16$\pm$0.43 43.86$\pm$0.82 45.01$\pm$0.71
HRDA 21.55$\pm$0.27 34.13$\pm$0.31 26.09$\pm$0.34 26.40$\pm$0.31 33.66$\pm$0.14 50.87$\pm$1.04 39.23$\pm$1.30 41.57$\pm$1.23
SeaFormer 20.34$\pm$0.16 25.33$\pm$0.21 25.03$\pm$0.13 24.26$\pm$0.16 27.85$\pm$0.64 35.80$\pm$0.55 32.58$\pm$0.58 33.48$\pm$0.56
TopFormer 20.41$\pm$0.00 25.19$\pm$0.45 25.20$\pm$0.22 24.28$\pm$0.20 29.84$\pm$1.73 38.09$\pm$2.72 34.29$\pm$1.75 35.41$\pm$2.08
@@
Methods Shift Dataset SynthiaSF Dataset
mIoU mPre mRec mF1 mIoU mPre mRec mF1
DeepLabv3+ 15.51$\pm$0.26 23.17$\pm$0.20 16.84$\pm$0.29 18.68$\pm$0.25 31.80$\pm$1.11 39.76$\pm$1.79 35.68$\pm$0.85 36.64$\pm$1.17
BiSeNetV2 10.74$\pm$2.50 15.23$\pm$5.01 11.88$\pm$2.81 12.63$\pm$3.38 27.04$\pm$0.68 33.49$\pm$0.93 31.42$\pm$0.64 31.49$\pm$0.70
SegNet 7.94$\pm$0.31 8.41$\pm$0.50 8.71$\pm$0.41 8.55$\pm$0.44 23.59$\pm$2.74 30.10$\pm$2.88 28.15$\pm$2.52 28.13$\pm$2.63
AttaNet 14.24$\pm$0.98 21.19$\pm$1.62 15.73$\pm$0.96 17.34$\pm$1.15 24.87$\pm$1.00 32.13$\pm$1.86 29.94$\pm$0.78 29.75$\pm$0.91
BASeg 16.60$\pm$0.18 22.93$\pm$0.30 17.87$\pm$0.18 19.91$\pm$0.18 33.70$\pm$0.54 54.05$\pm$0.44 36.61$\pm$0.57 40.48$\pm$0.47
HRDA 13.51$\pm$0.46 21.60$\pm$0.40 14.45$\pm$0.53 16.51$\pm$0.51 27.13$\pm$1.81 46.80$\pm$2.47 32.56$\pm$1.51 33.95$\pm$1.74
SeaFormer 7.69$\pm$0.08 8.26$\pm$0.03 8.43$\pm$0.09 8.33$\pm$0.05 26.01$\pm$0.87 33.23$\pm$0.69 30.30$\pm$0.76 30.64$\pm$0.71
TopFormer 13.53$\pm$1.52 20.00$\pm$2.83 15.01$\pm$1.69 16.44$\pm$2.04 25.45$\pm$0.76 32.35$\pm$0.60 30.20$\pm$0.58 30.31$\pm$0.63
AdvImmu (Ours) 26.09$\pm$1.54 29.35$\pm$2.06 26.54$\pm$1.20 26.47$\pm$1.21 51.30$\pm$4.36 74.91$\pm$2.46 52.19$\pm$4.01 55.65$\pm$3.79
Methods Cityscapes Dataset GTA5 Dataset
mIoU mPre mRec mF1 mIoU mPre mRec mF1
DeepLabv3+ 30.37$\pm$0.62 34.95$\pm$0.84 41.78$\pm$0.32 37.18$\pm$0.56 25.51$\pm$1.94 28.17$\pm$2.54 39.07$\pm$2.45 31.62$\pm$2.52
BiSeNetV2 19.33$\pm$1.20 20.61$\pm$1.66 25.47$\pm$1.30 22.34$\pm$1.52 16.33$\pm$0.43 17.29$\pm$0.43 28.10$\pm$0.26 20.30$\pm$0.42
SegNet 18.66$\pm$1.06 20.06$\pm$1.74 24.82$\pm$0.98 21.63$\pm$1.34 15.81$\pm$0.71 17.05$\pm$0.81 26.66$\pm$0.43 19.74$\pm$0.77
AttaNet 20.18$\pm$1.05 22.47$\pm$1.34 27.74$\pm$1.77 24.13$\pm$1.37 16.93$\pm$1.18 18.43$\pm$1.47 28.16$\pm$1.39 21.24$\pm$1.54
BASeg 64.56$\pm$0.60 82.29$\pm$0.76 73.15$\pm$0.62 76.44$\pm$0.60 58.09$\pm$0.76 79.08$\pm$0.57 64.42$\pm$0.61 68.72$\pm$0.66
HRDA 36.59$\pm$1.63 60.42$\pm$1.97 43.80$\pm$1.73 46.72$\pm$1.97 36.83$\pm$2.36 61.20$\pm$2.05 44.04$\pm$2.25 47.20$\pm$2.65
SeaFormer 22.53$\pm$0.42 25.59$\pm$0.45 30.19$\pm$0.39 26.90$\pm$0..38 17.84$\pm$1.22 19.35$\pm$1.48 29.54$\pm$1.55 22.21$\pm$1.58
TopFormer 20.81$\pm$1.06 23.07$\pm$1.32 28.50$\pm$1.77 24.88$\pm$1.46 19.33$\pm$1.42 21.47$\pm$1.81 31.29$\pm$1.75 24.64$\pm$1.91
AdvImmu (Ours) 71.14$\pm$3.70 88.64$\pm$4.12 79.69$\pm$3.84 79.17$\pm$2.88 58.11$\pm$2.47 78.89$\pm$2.16 72.15$\pm$2.00 67.58$\pm$2.36
TABLE II: Qualitative performance of AdvImmu against other SOTA baselines under various adverse weather conditions
        Raw Images Ground Truth         BASeg         HRDA DeepLabv3+ AdvImmu (Ours)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]